A study was conducted into the relationship between the age in years of a person and their average weekly after tax income.

You have been supplied with the following information regarding a regression model that had been developed as part of the study.

Regression analysis
sample intercept 646.58
sample slope 3.42

A particular observed value from the data is:

x = 41, y = 825.24

Calculate the residual of this observed value of the response variable. Give your answer to 2 decimal places.

residual = __________
Answer #1

y = 646.58 + 3.42 * x

Predicted value = 646.58 + 3.42 * 41 = 786.8

Observed value = 825.24

Residual = observed value - predicted value = 825.24 - 786.8 = 38.44 (ans)